# Magnetic Interaction in White Dwarf Binaries as Mechanism for Long-Period Radio Transients

### v1

@@ 5 Other sources
We have shown that the proposed magnetic interaction model invoking a WD – MD binary can successfully explain the observations of ILT J1101+5521. We suspect that the same model may apply to at least some of the other LPRTs as well, but probably with different parameters and observational geometric configurations. In the following, we discuss whether other reported LPRTs and related objects (Table 1) comply with the framework of our model.
  • Besides ILT J1101 + 5521 (de Ruiter et al., 2024), the most likely candidate that has the similar origin is the recently detected GLEAM-X J 0704-37 with a 2.9-hour period (Hurley-Walker et al., 2024). The polarization behavior of the source is variable with non-negligible linear (20–50%) and circular polarization (10–30%) over millisecond timescales. Optical observations revealed an MD counterpart (Hurley-Walker et al., 2024). Deeper observations will reveal whether the MD is in a binary system with a WD companion.
  • A very relevant system is the optical white dwarf pulsar AR Scorpii (Marsh et al., 2016). The optical emission can be understood within the framework of a binary interaction model invoking a bow shock between the WD and MD winds (Geng et al., 2016). It is likely that AR Scorpii and ILT J1101+5521 belong to very similar systems. The different apparent behaviors may be attributed to the longer period of AR Scorpii that gives a lower unipolar interaction power, and also possibly an unfavored geometry (e.g. a smaller inclination angle) that does not permit the radio emission to enter the line of sight.
@@
  • GLEAM-X J1627-52 with a period of 18.18 minutes (Hurley-Walker et al., 2022) and GPM J1839-10 with a period of 21 minutes (Hurley-Walker et al., 2023) are two relatively fast LPRTs. Their emissions have polarization properties more similar to pulsars, e.g. high linear polarization ($\Pi_{L}\sim 88\%$) for GLEAM-X J1627-52 (Hurley-Walker et al., 2022) and $\Pi_{L}\sim 10\%-100\%$ for GPM J1839-10 (Hurley-Walker et al., 2023), as well as flat polarization angle with occasional $90^{\circ}$ jumps. Deep searches for optical counterparts gave inconclusive results (Rea et al., 2022; Hurley-Walker et al., 2023). These two sources are least likely consistent with our WD – MD interaction model and may have a different origin.
  • ASKAPJ193505.1+214841.0 was reported to have a period of 53.8 minutes (Caleb et al., 2024). The polarization properties are highly variable: a bright pulse state with highly linear polarization ($>90\%$) and a weak pulse state ($\sim$26 times fainter than the bright state) with highly circular polarization. There is a near IR source in the field, but whether it is associated with the radio emitter is uncertain. Further observations are needed to establish whether the IR source is an MD associated with the LPRT, and whether the source could be understood within the framework of the WD – MD interaction model.
Table 1: Published LPRTs with measured periods, luminosities, polarization degrees and possible optical counterparts. AR Scorpii is also included because of its possible similar origin.
Source $P\ ({\rm min})$ $\Pi_{L}\ (\%)$ Optical counterparts Reference
GCRT J1745–3009 77 / No detection Hyman et al. (2005); Kaplan et al. (2008)
GLEAM-X J1627-52 18.18 $88\pm 1$ No detection Rea et al. (2022); Hurley-Walker et al. (2022)
GPM J1839-10 21 10–100 KD or MD? Hurley-Walker et al. (2023)
ASKAPJ193505.1+214841.0 53.8 $>90$ near IR source ? Caleb et al. (2024)
ILT J1101 + 5521 125.5 $51\pm 6$ WD+MD de Ruiter et al. (2024)
GLEAM-X J 0704-37 174 20–50 MD Hurley-Walker et al. (2024)
AR Scorpii 213.6 / WD+MD Marsh et al. (2016)
@@ 6 Conclusions
In this paper, we propose a magnetic interaction model invoking a magnetized WD and a non-magnetized MD through the unipolar induction effect to interpret the newly discovered LPRT, ILT J1101+5521. We argue that a relativistic version of the electron cyclotron maser emission is the most likely emission mechanism. This is a more energetic version of the Jupiter-Io interaction observed in the solar system.
We further explore the possibility of applying this model to the emergent LPRT population with a rapidly growing number. We find that those with relative long periods such as GLEAM-X J0704-37 and GCRT J1745-3009 and probably ASKAPJ193505.1+214841.0 as well, may be interpreted within the same theoretical framework. The WD optical pulsar AR Scorpii may also belong to the same type of system with a different evolution stage or viewing geometry. However, we note that the sources with relatively shorter periods, such as GLEAM-X J1627-52 and GPM J1839-10, do not seem to meet the model predictions and may have a different physical origin, e.g. a single WD pulsar or a long-period magnetar. More multi-wavelength observations of these systems and new sources will help to test the suggested dichotomy and the unified picture of the ultra-long period sub-population of LPRTs.
@@ Acknowledgements
This work is supported by the Nevada Center for Astrophysics, NASA 80NSSC23M0104 and a Top Tier Doctoral Graduate Research Assistantship (TTDGRA) at University of Nevada, Las Vegas.

### v2

@@ 5 Other sources
We have shown that the proposed magnetic interaction model invoking a WD – MD binary can successfully explain the observations of ILT J1101+5521. We suspect that the same model may apply to at least some of the other LPRTs as well, but probably with different parameters and observational geometric configurations. In the following, we discuss whether other reported LPRTs and related objects (summarized in Table 1) align with the framework of our model:
  • Besides ILT J1101 + 5521 (de Ruiter et al., 2024), the most likely candidate that has the similar origin is the recently detected GLEAM-X J 0704-37 with a 2.9-hour period (Hurley-Walker et al., 2024). The polarization behavior of the source is variable with non-negligible linear (20–50%) and circular polarization (10–30%) over millisecond timescales. Optical observations revealed an MD counterpart (Hurley-Walker et al., 2024). Deeper observations will reveal whether the MD is in a binary system with a WD companion. The observed mixed linear and circular polarization can be produced when the line of sight is inside the $1/\gamma$ cone in this model.
  • A very relevant system is the optical white dwarf pulsar AR Scorpii with a period of 213.6 minutes (Marsh et al., 2016; Buckley et al., 2017). The observed broadband spectra of AR Scorpii at high frequency band is consistent with synchrotron radiation via relativistic electrons. The polarization properties of AR Scorpii is quite different from other intermediate polars with dominant circular polarization which might be produced through cyclotron emission. The optical observations of AR Scorpii show that the linear polarization ($\Pi_{L}\sim 40\%$) and both low circular polarization ($\Pi_{V}$ is a few percent) are consistent with synchrotron emission (Buckley et al., 2017), probably originates from a bow shock at the interface of the WD and MD winds (Geng et al., 2016). Radio emission from AR Scorpii seems to have two origins. Above 10 GHz, the radio emission is not polarized and it is consistent with self-absorbed synchrotron radiation. Below 10 GHz, it shows weak linear polarization and a circular polarization with $\Pi_{V}\sim 30\%$. Its pulsation follows the beat frequency instead of the orbital frequency, which is different from ILT J1101+5521. It was suggested that this emission originates from non-relativistic ECME (Stanway et al., 2018). Within our framework, the beamed relativistic ECME emission in the current loop region might be missed due to a unfavorable geometry, and the observed radio emission may be related to reprocessed emission from the a non-relativistic plasma outside of the current loop region.
@@
  • GLEAM-X J1627-52 with a period of 18.18 minutes (Hurley-Walker et al., 2022) and GPM J1839-10 with a period of 21 minutes (Hurley-Walker et al., 2023; Men et al., 2025) are two relatively fast LPRTs. Their emissions have polarization properties more similar to pulsars, e.g. high linear polarization ($\Pi_{L}\sim 88\%$) for GLEAM-X J1627-52 (Hurley-Walker et al., 2022) and $\Pi_{L}\sim 10\%-100\%$ for GPM J1839-10 (Hurley-Walker et al., 2023), as well as flat polarization angle with occasional $90^{\circ}$ jumps. Deep searches for optical counterparts gave inconclusive results (Rea et al., 2022; Hurley-Walker et al., 2023). These two sources are least likely consistent with our WD – RD interaction model and may have a different origin. However, if these sources are interpreted within our model, the observed high linear polarization requires that the line of sight is close to the edge of the radiation cone ($\theta_{v}\sim 1/\gamma$).
  • ASKAPJ193505.1+214841.0 was reported to have a period of 53.8 minutes (Caleb et al., 2024). The polarization properties are highly variable: a bright pulse state with highly linear polarization ($>90\%$) and a weak pulse state ($\sim$26 times fainter than the bright state) with highly circular polarization. Within this model, the two emission components correspond to two emission regions with different viewing angles, with the bright component observed near the radiation cone edge and the fainter component observed within the radiation cone. There is a near IR source in the field, but whether it is associated with the radio emitter is uncertain. Further observations are needed to establish whether the IR source is an RD associated with the LPRT, and whether the source could be understood within the framework of the WD – RD interaction model.
  • DART J1832-0911 and ASKAP J1832-0911 are the same LPRT source with a 44.2-minute period detected by DAocheng Radio Telescope (DART) and Australian SKA Pathfinder (ASKAP), respectively (Li et al., 2024; Wang et al., 2024). The source was found to be associated with a supernova remnant, favoring a young neutron star origin (Li et al., 2024). The radio emission is $\sim 100\%$ linear polarized with a relatively flat polarization angle across the pulses and the frequency bandwidth is relatively broad. (Li et al., 2024). The radio emission polarization properties are also reported to be highly linear polarized with a total polarization degree $\Pi_{p}\simeq 92\pm 3\%$. Linear polarization is $\Pi_{L}\sim 75\%$ and circular polarization is $\Pi_{V}\sim 50\%$ (Wang et al., 2024). The source also has an X-ray emission counterpart with luminosity $\sim 10^{33}\ {\rm erg\ s^{-1}}$ (Wang et al., 2024). This source is very likely related to a newborn neutron star, not related to the mechanism discussed in this paper.
  • ASKAP J183950.5-075635.0 was reported to have a period of 387 minutes (Lee et al., 2025). The emission is highly polarized with a linear polarization degree $\sim 90\%$ and a circular polarization degree $\sim 37\%$. This source can be interpreted within the framework of our model, with the line of sight slightly within the $1/\gamma$ cone. The narrow spectrum is also consistent with the prediction of the relativistic ECME model.
@@ 6 Conclusions
In this paper, we propose a magnetic interaction model invoking a magnetized WD and a weakly magnetized RD through the unipolar induction effect to interpret the newly discovered LPRT, ILT J1101+5521. We argue that a relativistic version of the ECME is the most likely emission mechanism. This is a more energetic version of the Jupiter-Io interaction observed in the solar system.
We further explore the possibility of applying this model to the emergent LPRT population with a rapidly growing number. We find that those with relative long periods such as GLEAM-X J0704-37 and GCRT J1745-3009 and probably ASKAPJ193505.1+214841.0 and ASKAP J183950.5-075635.0 as well, may be interpreted within the same theoretical framework. The WD optical pulsar AR Scorpii may also belong to the same type of system with a different viewing geometry. The asynchronous configuration indicates large electric potential and relativistic particles. Relativistic ECME can produce both high linear polarization (if the line of sight is close to the $1/\gamma$ cone) and high circular polarization (if the line of sight is well within the $1/\gamma$ cone).
On the other hand, we note that the sources with relatively shorter periods, such as GLEAM-X J1627-52, GPM J1839-10, and DART/ASKAP J1832-0911, may have a different physical origin, e.g. a single WD pulsar or a long-period magnetar. More multi-wavelength observations of these systems and new sources will help to test the suggested dichotomy and the unified picture of the ultra-long period sub-population of LPRTs. The supernova remnant association with DART/ASKAP J1832-0911 (Li et al., 2024) provides the first observational clue of such dichotomy.
We also predict high-energy emissions from LPRTs within the framework of our model. X-rays can be produced via relativistically boosted cyclotron radiation with the maximum Lorentz factor and both X-rays and $\gamma$-rays can be produced via IC scattering of relativistic particles off the thermal photons from WD surface. We find that the observed X-ray counterpart of AR Scorpii can be achieved in the WD-RD system, but that of and DART/ASKAP J1832-0911 may originate from other systems, possibly invoking a newborn magnetar. In the future, the detection of X-ray counterparts and possibly $\gamma$-ray counterparts could be used to distinguish different engines of LPRTs.
@@ Acknowledgements
We thank Manisha Caleb, Di Li, Yunpeng Men, Nanda Rea, Kaustubh Rajwade, and Ziteng Wang for discussion and an anonymous referee for very helpful comments. This work is supported by the Nevada Center for Astrophysics and NASA 80NSSC23M0104.
Appendix A Polarization of relativistic ECME
In this appendix, we provide a calculation of the polarization properties of relativistic ECME. We first treat the problem in the comoving frame of the relativistic electron that moves along background magnetic field with Lorentz factor $\gamma$, and then perform Lorentz transformation to study radiation back to the lab frame. In the comoving frame, the electron gyrates around the background magnetic field and experiences cyclotron or synchrotron radiation.
Consider an electron gyrates around the background magnetic field non-relativistically and emits cyclotron radiation. We define the viewing angle $\theta_{v}^{\prime}$ between the line of sight and background magnetic field. The circular and linear polarization degrees can be written as
$\Pi_{V,\rm cyc}=-\frac{2\cos\theta_{v}^{\prime}}{1+\cos^{2}\theta_{v}^{\prime}}\ \ \&\ \ \Pi_{L,\rm cyc}=\frac{1-\cos^{2}\theta_{v}^{\prime}}{1+\cos^{2}\theta_{v}^{\prime}}.$ (A1)
One can see that when the line of sight is parallel to the gyration plane, i.e. $\theta_{v}^{\prime}=\pi/2$, the radiation is 100% linearly polarized in the comoving frame. When the line of sight is along the background magnetic field, i.e. $\theta_{v}^{\prime}=0$, the radiation is 100% circularly polarized.
Now consider emission in the lab frame. The Lorentz transformation on the viewing angle is
$\sin\theta_{v}^{\prime}={\cal D}\sin\theta_{v},\ {\cal D}=\frac{1}{\gamma(1-\beta\cos\theta_{v})},$ (A2)
where ${\cal D}$ is the Doppler factor, $\gamma$ is the Lorentz factor of the electron moving along the background magnetic field. When $\gamma\gg 1$, $\theta_{v}^{\prime}=\pi/2$ corresponds to $\theta_{v}\sim 1/\gamma$. Therefore, when the line of sight is along $\theta_{v}\sim 1/\gamma$, the observer will detect $\sim 100\%$ linear polarization. When $\theta_{v}=0$, the observer will detect $\sim 100\%$ circular polarization. More generally, the radiation is elliptically polarized in the range of viewing angle $0<\theta_{v}<1/\gamma$. Since the chance of $\theta_{v}=0$ is negligibly small, one does not expect purely circularly polarized emission for relativistic ECME. Since the line of sight probability $p(\theta_{v})d\theta_{v}=\sin\theta_{v}d\theta_{v}$, one expects that most viewing angles carry significant linear polarization. Outside the $1/\gamma$ cone, the radiation intensity drops rapidly, the chance of observing an object in these viewing angles diminishes.